OpenStudy (anonymous):

Integrate the following:

OpenStudy (anonymous):

\[∫ \frac{ 5 }{ \sqrt{4-x ^{2}} }dx\]

OpenStudy (hba):

Did ya try solvin it ?

OpenStudy (hba):

@bmelyk

OpenStudy (anonymous):

i dont know how to state it.

OpenStudy (anonymous):

start*

OpenStudy (slaaibak):

hint: use the substitution x = 2sin u

OpenStudy (anonymous):

how do i start it though?

OpenStudy (slaaibak):

using x = 2sin u dx = 2cos u \[\int\limits {5 \times 2 \cos u \over \sqrt{ 4 - 4\sin^2 u}} du = \int\limits { 5 \times 2\cos u \over 2\sqrt{1-\sin^2u}}du\] Try taking it from here

OpenStudy (anonymous):

i dont get what you are doing there.

OpenStudy (slaaibak):

tell me which part you don't understand? The choice of substitution, the substitution itself?

OpenStudy (anonymous):

choice of substitution